OpenStudy (study312):

Allison has a poster that is 15 in by 18 in. What will the dimensions of the poster be if she scales it down by a factor of one-third ? 15 in by 18 in 45 in by 54 in 2.5 in by 2 in 5 in by 6 in

OpenStudy (michele_laino):

hint: if we scale down the original dimensions by 1/3, then the new dimensions will be these ones: \[\Large \begin{gathered} 15 \to \frac{{15}}{3} = ...? \hfill \\ \hfill \\ 18 \to \frac{{18}}{3} = ...? \hfill \\ \end{gathered} \]

OpenStudy (study312):

so 15 ÷ 3 and 18 ÷ 3?

OpenStudy (michele_laino):

yes!

OpenStudy (study312):

oh ok

OpenStudy (study312):

so 15 ÷ 3 = 5 and 18 ÷ 3 = 6. Ok thx I get it now, I was really confused because I was doing 1 ÷ 3 • 15 and 18.

OpenStudy (michele_laino):

ok! :)